I. I want to spend no less than $3,000 and no more than $7000 on my landscaping project.
a) Write an inequality that shows how much money I can spend. Let “m” represent the money. (3 pts)

b) For the first phase of my project I want to spread decorative rocks and plant some hibiscus bushes. I need 35 tons of rock to cover the area. If each ton costs $50 and each bush costs $25 how many bushes can I buy with a budget of $2000? Write an inequality that illustrates the problem then solve it (let “h” represent the hibiscus bushes). Express your answer as an inequality and explain how you arrived at your answer. (3 pts)
c) Can I buy 7 hibiscus bushes? Explain why or why not.

Answer by rfer(16322)

A) 3000 <= m <= 7000
B)
35*50=$1750
2000-1750=$250
250/25=10 plants
C) you should be able to answer